Consider the following examples in real word situations that describes a correspondence between phenomena

The total cost for repairing a cell phone is R320 for parts plus R50 per hour for labour.

Siphokazi sells 70% alcohol based hand sanitizers.she earns R230 a week and receives a commission of 15% of her sales.

A rental van costs R230 per day plus R4,30 per kilometre.


1
Expert's answer
2021-07-02T07:24:49-0400


1: Total cost=cost of parts+cost per lab
320+50=R370

2: Total weekly earning= salary+commission

   Commision=15/100x230=34.5
Total weekly earning=34.50+230
=R264.50

3: Total amount of van rental=cost per day+cost per kilometer

"230+430=R660"